# Rectified Flows for Fast Multiscale Fluid Flow Modeling

### v1

@@ D.2 Training Efficiency Observations
$u_{x}$ Mean Error ($e_{\mu}$) 280,000 3.33 ×$\times 10^{-2}$ 3.45 ×$\times 10^{-2}$
$u_{y}$ Mean Error ($e_{\mu}$) 400,000 1.70 ×$\times 10^{-1}$ 1.88 ×$\times 10^{-1}$
$u_{x}$ Std Error ($e_{\sigma}$) 400,000 6.48 ×$\times 10^{-2}$ 7.24 ×$\times 10^{-2}$
$u_{y}$ Std Error ($e_{\sigma}$) 340,000 7.91 ×$\times 10^{-2}$ 7.97 ×$\times 10^{-2}$
$u_{x}$ Wasserstein-1 ($W_{1}$) 200,000 3.89 ×$\times 10^{-2}$ 3.95 ×$\times 10^{-2}$
$u_{y}$ Wasserstein-1 ($W_{1}$) 4.11 ×$\times 10^{-2}$
Figure D.1: Training efficiency comparison. Performance of ReFlow and GenCFD on the Macro-Micro SL ensemble dataset evaluated at multiple training checkpoints. Metrics match those used in the main text: relative $L^{2}$ error of the mean ($e_{\mu}$) and standard deviation ($e_{\sigma}$), and the average 1-point Wasserstein-1 distance (W1), computed for both velocity components ($u_{x}$, $u_{y}$). The $y$-axis uses a logarithmic scale, while the $x$-axis (training iteration) is linear.
D.3 Experimental Advantages of Straightness
@@
  • Latent-Space Trajectory Visualization via PCA via PCA.
  • Per-Sample Average Error Evolution: average MSE across physical channels.
  • Evolution of the Energy Spectrum: radial power-spectrum of energy $E$ as inpaintig progresses.
@@
Figure D.2: Density trajectories for Sample 1: ReFlow (top) vs. GenCFD (bottom). ReFlow already recovers sharp shock fronts by diffusion time $t=1.0$.
Figure D.3: Density trajectories for Sample 2.
Figure D.4: Density trajectories for Sample 3.
Latent-Space Trajectory Visualization via PCA
@@
Figure D.5: 3D PCA latent-space trajectories (Samples 1–6). We project the flattened snapshots at each of $T$ timesteps into the top three PCA components. Rectified Flow (blue) consistently traces a straighter path than GenCFD (red), confirming its “rectified” evolution. Circles mark $t=0$, squares mark $t=1$, and intermediate timesteps are annotated.
To further illustrate the nature of each model’s generative dynamics, we visualize single-sample generation trajectories from both models in a 2D PCA space for the RM dataset (Figure D.6). For each physical channel ($\rho$, $m_{x}$, $m_{y}$, $p$), we apply PCA to the full set of ground truth snapshots to define a common projection basis. We then overlay the generative trajectory of a single sample in this space. To contextualize the generated dynamics, we add kernel density estimates (KDEs) showing the empirical data distribution in the same projected space.
@@
Figure D.6: 2D PCA latent-space trajectories. We project the flattened velocity snapshots from the RM dataset into a 2D PCA space, computed independently for each channel ($\rho$, $m_{x}$, $m_{y}$, $p$) across all $T$ time steps. To visualize the underlying distribution of target fields, we apply kernel density estimation (KDE) per channel. The ReFlow model (top) evolves over 10 rectified time steps, producing visibly straighter and more structured trajectories compared to the score-based diffusion model GenCFD (bottom), which uses 128 stochastic steps. Although the projection into two principal components necessarily truncates the high-dimensional dynamics, the qualitative difference remains evident, illustrating the smoother and more directed progression learned by ReFlow.
Per-Sample Average Error Evolution
We compute the mean-squared error at each timestep $t$ and average across all $C$ channels:
@@
Figure D.7 shows, for Samples 1–6, the log-scaled average MSE (left) and the normalized error $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$ (right). Rectified Flow (blue) consistently reduces error faster and to a lower residual than GenCFD (red).
@@
Figure D.7: Per-sample average MSE evolution. Each subfigure plots, left half: $\log_{10}\overline{\mathrm{MSE}}(t)$; right half: $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$. Rectified Flow (blue) rapidly drives error down and keeps it orders of magnitude below GenCFD (red) across all samples.
@@
Figure D.8: Energy spectrum evolution for Sample 1. ReFlow (blue) outperforms GenCFD (red) in capturing high-wavenumber energy.
Figure D.9: Energy spectrum evolution for Sample 2. ReFlow (blue) remains closer to the ground truth (dashed) than GenCFD (red).
Figure D.10: Energy spectrum evolution for Sample 3. The inset log-MSE annotations quantify ReFlow’s advantage at each $t$.
Figure D.11: Energy spectrum evolution for Sample 4. ReFlow’s spectrum (blue) gets closer to the target (dashed) as $\tau\to 1.0$.
Figure D.12: Energy spectrum evolution for Sample 5. ReFlow’s superior high-$k$ fidelity is evident throughout.
Figure D.13: Energy spectrum evolution for Sample 6. ReFlow better captures the ground-truth spectrum’s slope across all timesteps.
Appendix E Computational Considerations
[…]
Table E.1: Aggregate hardware utilization for all “usual-setup” runs.
@@ Appendix E Computational Considerations
Since all datasets are paired with models of identical size and the hardware envelope (GPU power, memory, temperature, CPU threads, I/O) is effectively the same across runs, any observed performance differences are attributable to the sampling strategy (noise schedule, ODE steps) rather than to model-size or hardware variations (the network traffic is due to Wandb usage).
@@
Appendix F Additional Visualizations
Test-time ensemble perturbation (evaluation only).
@@ Appendix F Additional Visualizations
Following the protocol of GenCFD [21], we construct small “micro-ensembles” around each test initial condition $\bar{u}$ by sampling $M_{\mathrm{micro}}$ perturbed fields uniformly in an $\varepsilon$–ball centered at $\bar{u}$. Each perturbed member is then advanced to the evaluation time $t$ with a high-fidelity PDE solver to produce a reference ensemble $\{u^{(j)}_{\rm ref}(t)\}_{j=1}^{M_{\mathrm{micro}}}$. Our model generates a corresponding ensemble $\{u^{(j)}_{\rm model}(t)\}_{j=1}^{M_{\mathrm{micro}}}$ from the same perturbed inputs, but without ever observing ensembles during training (which only uses independent pairs $(u_{0},u_{1})$). We compare these two ensembles via mean-field error $e_{\mu}$, standard-deviation error $e_{\sigma}$, and spatially averaged 1-Wasserstein distance $\overline{W}_{1}$ (see SM §D).
@@
Out-of-distribution testing (cloud–shock and shear-flow). Because the micro-perturbations render each initial field slightly OOD, this evaluation probes the model’s ability to generalize under small chaotic perturbations. In these tasks, our Rectified Flow ensemble closely tracks the reference uncertainty and multiscale statistics, whereas deterministic baselines collapse to a single trajectory and fail to capture the conditional variability.
@@
In-distribution scaling (Richtmyer–Meshkov). For the Richtmyer–Meshkov dataset, we additionally study the effect of training set size using a held-out test set from the same distribution. Here, even deterministic approaches improve with more data, but still underperform on the micro-ensemble test: they require significantly larger training sets to match ReFlow’s accuracy and exhibit poor calibration under small perturbations.
@@
In this appendix we provide a comprehensive set of qualitative comparisons between the baselines (GenCFD variants, UViT, FNO) models and our Rectified Flow (ReFlow) approach across three representative 2D fluid-flow benchmarks (Richtmyer–Meshkov, shear layer roll-up, and cloud-shock interaction). For each task we show:
@@
  • Mean fields Ensemble averages reveal how each deterministic model diffuses fine-scale features (UVit/FNO) versus perserving sharp interfaces (ReFlow, GenCFD variants).
@@
  • Uncertainty maps Per-pixel standard deviations illustrate whether uncertainty localizes in physically meaningful regions (e.g. shocks, shear interfaces) or remains overly diffuse.
  • Random samples To qualitatively assess generative fidelity, we include representative random samples from ReFlow alongside outputs from selected baselines, specifically the deterministic FNO and the hybrid GenCFD $\boldsymbol{\circ}$ FNO. The latter combines FNO’s low-frequency structural prediction with GenCFD’s generative modeling of fine-scale turbulence. This comparison highlights ReFlow’s ability to produce realistic, high-resolution flow fields that preserve physical coherence across different scales.
@@
  • Spectral & pointwise comparisons (Fig. F.4: Richtmyer–Meshkov global $E(k)$, local spectra, and pointwise density histograms; Fig. F.7: Shear-layer global $u_{x}$ spectra and pointwise $u_{x}$ distributions; Fig. F.11: Cloud–shock global $m_{y}$ spectra and pointwise energy distributions) — combined Fourier-space and histogram panels quantify multiscale fidelity and calibration of the predictive ensembles.
@@
Together, these visualizations underscore ReFlow’s ability to maintain sharper physical structures, produce better-calibrated uncertainties, and faithfully reproduce multiscale spectral statistics.
@@
Figure F.1: Richtmyer–Meshkov Mean Field Comparison: Mean velocity fields for the Richtmyer–Meshkov dataset, shown across all physical channels and models. In many cases, predictions closely resemble the ground truth, with differences often imperceptible, underscoring the overall high quality and consistency of the generated mean fields.
@@
Figure F.2: Richtmyer–Meshkov Uncertainty Comparison: Per-pixel standard deviation fields from the Richtmyer–Meshkov dataset across all models. Stochastic generative models capture spatially coherent uncertainty patterns, outperforming the deterministic baselines (FNO and UViT), which exhibit limited variability and poor uncertainty localization.
Figure F.3: Richtmyer–Meshkov Individual Sample Comparison: A representative random test sample from the Richtmyer–Meshkov dataset, visualized across all four physical channels ($\rho$, $m_{x}$, $m_{y}$, $p$). All models produce physically plausible results, with only little variation in fidelity to fine-scale structures. ReFlow offers the closest visual match to the ground truth, preserving coherent details and sharp interfaces. Notably, the GenCFD $\boldsymbol{\circ}$ FNO hybrid enhances spectral richness by reconstructing high-frequency features based on a low-frequency FNO prior, illustrating its capacity to recover sharper modes even when starting from a coarse prediction.
@@
Figure F.4: Richtmyer–Meshkov: spectral and pointwise distribution comparisons between UVit and Rectified Flow.
@@
Figure F.5: Shear-Layer Mean and Standard Deviation Comparison: Visualization of mean-field ($\mu$) and standard-deviation ($\sigma$) predictions for the Shear-Layer dataset across both velocity components ($u_{x}$, $u_{y}$). For each model, the left half shows the predicted mean fields and the right half shows per-pixel standard deviations. The top row contains the reference ground truth. Despite the reduced channel complexity of this dataset, clear differences emerge in how well models capture coherent structures and localized uncertainties. ReFlow preserves sharper features and more accurately reflects the spatial distribution of uncertainty.
Figure F.6: Shear-Layer Individual Sample Comparison. A representative random test sample from the Shear-Layer dataset, shown across both physical channels ($u_{x},u_{y}$). All models generate physically plausible results, though with varying degrees of detail and sharpness. ReFlow and GenCFD $\boldsymbol{\circ}$FNO most closely resemble the ground truth, preserving fine-scale structures and coherent interfaces. GenCFD $\boldsymbol{\circ}$FNO enhances high-frequency detail by leveraging the low-frequency prior from the initial FNO prediction.
Figure F.7: Shear-layer: global $u_{x}$ spectra and pointwise $u_{x}$ distributions.
@@
Figure F.8: Cloud-Shock Mean Field Comparison: Mean velocity fields for the Cloud-Shock dataset, shown across all physical channels and models.
@@
Figure F.9: Cloud–Shock Uncertainty Comparison: Per-pixel standard deviation visualizations for the Cloud–Shock dataset across all models.
Figure F.10: Cloud–Shock Individual Sample Comparison: A representative random test sample from the Cloud–Shock dataset, displayed across all physical channels. While all models produce plausible results, ReFlow and the hybrid GenCFD $\boldsymbol{\circ}$ FNO most closely align with the ground truth, particularly in preserving fine-scale detail. The spectral reconstruction effect of the hybrid model is especially evident, recovering high-frequency structures from the coarse FNO prediction and enhancing the overall visual fidelity.
Figure F.11: Cloud-shock: spectral and pointwise energy comparisons.

### v2

@@ Appendix I From law-level bounds to reported pointwise metrics
$u_{x}$ Mean Error ($e_{\mu}$) 280,000 3.33 $\times 10^{-2}$ 3.45 $\times 10^{-2}$
$u_{y}$ Mean Error ($e_{\mu}$) 400,000 1.70 $\times 10^{-1}$ 1.88 $\times 10^{-1}$
$u_{x}$ Std Error ($e_{\sigma}$) 400,000 6.48 $\times 10^{-2}$ 7.24 $\times 10^{-2}$
$u_{y}$ Std Error ($e_{\sigma}$) 340,000 7.91 $\times 10^{-2}$ 7.97 $\times 10^{-2}$
$u_{x}$ Wasserstein-1 ($W_{1}$) 200,000 3.89 $\times 10^{-2}$ 3.95 $\times 10^{-2}$
$u_{y}$ Wasserstein-1 ($W_{1}$) 4.11 $\times 10^{-2}$
Figure 6: Training efficiency comparison. Performance of ReFlow and GenCFD on the Macro-Micro SL ensemble dataset evaluated at multiple training checkpoints. Metrics match those used in the main text: relative $L^{2}$ error of the mean ($e_{\mu}$) and standard deviation ($e_{\sigma}$), and the average 1-point Wasserstein-1 distance (W1), computed for both velocity components ($u_{x}$, $u_{y}$). The $y$-axis uses a logarithmic scale, while the $x$-axis (training iteration) is linear.
M.3 Experimental Advantages of Straightness
@@
  • Latent‐Space Trajectory Visualization via PCA via PCA.
  • Per‐Sample Average Error Evolution: average MSE across physical channels.
  • Evolution of the Energy Spectrum: radial power‐spectrum of energy $E$ as inpaintig progresses.
@@
Figure 7: Density trajectories for Sample 1: ReFlow (top) vs. GenCFD (bottom). ReFlow already recovers sharp shock fronts by diffusion time $t=1.0$.
Figure 8: Density trajectories for Sample 2.
Figure 9: Density trajectories for Sample 3.
Latent‐Space Trajectory Visualization via PCA
@@
Figure 10: 3D PCA latent-space trajectories (Samples 1–6). We project the flattened snapshots at each of $T$ timesteps into the top three PCA components. Rectified Flow (blue) consistently traces a straighter path than GenCFD (red), confirming its “rectified” evolution. Circles mark $t=0$, squares mark $t=1$, and intermediate timesteps are annotated.
To further illustrate the nature of each model’s generative dynamics, we visualize single-sample generation trajectories from both models in a 2D PCA space for the RM dataset (Figure 11). For each physical channel ($\rho$, $m_{x}$, $m_{y}$, $p$), we apply PCA to the full set of ground truth snapshots to define a common projection basis. We then overlay the generative trajectory of a single sample in this space. To contextualize the generated dynamics, we add kernel density estimates (KDEs) showing the empirical data distribution in the same projected space.
@@
Figure 11: 2D PCA latent-space trajectories. We project the flattened velocity snapshots from the RM dataset into a 2D PCA space, computed independently for each channel ($\rho$, $m_{x}$, $m_{y}$, $p$) across all $T$ time steps. To visualize the underlying distribution of target fields, we apply kernel density estimation (KDE) per channel. The ReFlow model (top) evolves over 10 rectified time steps, producing visibly straighter and more structured trajectories compared to the score-based diffusion model GenCFD (bottom), which uses 128 stochastic steps. Although the projection into two principal components necessarily truncates the high-dimensional dynamics, the qualitative difference remains evident, illustrating the smoother and more directed progression learned by ReFlow.
Per‐Sample Average Error Evolution
We compute the mean‐squared error at each timestep $t$ and average across all $C$ channels:
@@
Figure 12 shows, for Samples 1–6, the log‐scaled average MSE (left) and the normalized error $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$ (right). Rectified Flow (blue) consistently reduces error faster and to a lower residual than GenCFD (red).
@@
Figure 12: Per‐sample average MSE evolution. Each subfigure plots, left half: $\log_{10}\overline{\mathrm{MSE}}(t)$; right half: $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$. Rectified Flow (blue) rapidly drives error down and keeps it orders of magnitude below GenCFD (red) across all samples.
@@
Figure 13: Energy spectrum evolution for Sample 1. ReFlow (blue) outperforms GenCFD (red) in capturing high-wavenumber energy.
Figure 14: Energy spectrum evolution for Sample 2. ReFlow (blue) remains closer to the ground truth (dashed) than GenCFD (red).
Figure 15: Energy spectrum evolution for Sample 3. The inset log-MSE annotations quantify ReFlow’s advantage at each $t$.
Figure 16: Energy spectrum evolution for Sample 4. ReFlow’s spectrum (blue) gets closer to the target (dashed) as $\tau\to 1.0$.
Figure 17: Energy spectrum evolution for Sample 5. ReFlow’s superior high-$k$ fidelity is evident throughout.
Figure 18: Energy spectrum evolution for Sample 6. ReFlow better captures the ground‐truth spectrum’s slope across all timesteps.
Appendix N Computational Considerations
All of our ReFlow models for 2D fluid-flow tasks use the same network configuration summarized in Table 5 (channels $(128,256,256)$, downsample ratios $(2,2,2)$, 8 attention heads, dropout 0.1, batch size 16, etc.). Consequently, every model—regardless of dataset—has the same total parameter count (approximately $8\text{–}10$M parameters) and identical layer‐wise dimensions. We ran each training job for 100 000 iterations on a single NVIDIA Quadro RTX 6000 (Driver 570.124.06, CUDA 12.8) in persistence mode. Resource utilization averaged as follows over the full 100 h window:
Table 8: Aggregate hardware utilization for all “usual‐setup” runs.
@@
Since all datasets are paired with models of identical size and the hardware envelope (GPU power, memory, temperature, CPU threads, I/O) is effectively the same across runs, any observed performance differences are attributable to the sampling strategy (noise schedule, ODE steps) rather than to model‐size or hardware variations (the network traffic is due to Wandb usage).
@@
Appendix O Code and Data Availability
Prior to acceptance, we will make available all three datasets, the complete data-processing scripts, and a fully functional training and evaluation pipeline for ReFlow, including a provided model checkpoint that can be loaded and tested. A more comprehensive, user-friendly code release will follow shortly thereafter.
Appendix P Additional Visualizations
Test‐time ensemble perturbation (evaluation only).
@@
Following the protocol of GenCFD Molinaro et al. (2024), we construct small “micro‐ensembles” around each test initial condition $\bar{u}$ by sampling $M_{\mathrm{micro}}$ perturbed fields uniformly in an $\varepsilon$–ball centered at $\bar{u}$. Each perturbed member is then advanced to the evaluation time $t$ with a high‐fidelity PDE solver to produce a reference ensemble $\{u^{(j)}_{\rm ref}(t)\}_{j=1}^{M_{\mathrm{micro}}}$. Our model generates a corresponding ensemble $\{u^{(j)}_{\rm model}(t)\}_{j=1}^{M_{\mathrm{micro}}}$ from the same perturbed inputs, but without ever observing ensembles during training (which only uses independent pairs $(u_{0},u_{1})$). We compare these two ensembles via mean‐field error $e_{\mu}$, standard‐deviation error $e_{\sigma}$, and spatially averaged 1‐Wasserstein distance $\overline{W}_{1}$ (see SM §M).
@@
Out‐of‐distribution testing (cloud–shock and shear‐flow). Because the micro‐perturbations render each initial field slightly OOD, this evaluation probes the model’s ability to generalize under small chaotic perturbations. In these tasks, our Rectified Flow ensemble closely tracks the reference uncertainty and multiscale statistics, whereas deterministic baselines collapse to a single trajectory and fail to capture the conditional variability.
@@
In‐distribution scaling (Richtmyer–Meshkov). For the Richtmyer–Meshkov dataset, we additionally study the effect of training set size using a held‐out test set from the same distribution. Here, even deterministic approaches improve with more data, but still underperform on the micro‐ensemble test: they require significantly larger training sets to match ReFlow’s accuracy and exhibit poor calibration under small perturbations.
@@
In this appendix we provide a comprehensive set of qualitative comparisons between the baselines (GenCFD variants, UViT, FNO) models and our Rectified Flow (ReFlow) approach across three representative 2D fluid‐flow benchmarks (Richtmyer–Meshkov, shear layer roll‐up, and cloud-shock interaction). For each task we show:
@@
  • Mean fields Ensemble averages reveal how each deterministic model diffuses fine‐scale features (UVit/FNO) versus perserving sharp interfaces (ReFlow, GenCFD variants).
@@
  • Uncertainty maps Per‐pixel standard deviations illustrate whether uncertainty localizes in physically meaningful regions (e.g. shocks, shear interfaces) or remains overly diffuse.
[…]
@@
  • Spectral & pointwise comparisons (Fig. 22: Richtmyer–Meshkov global $E(k)$, local spectra, and pointwise density histograms; Fig. 25: Shear‐layer global $u_{x}$ spectra and pointwise $u_{x}$ distributions; Fig. 29: Cloud–shock global $m_{y}$ spectra and pointwise energy distributions) — combined Fourier‐space and histogram panels quantify multiscale fidelity and calibration of the predictive ensembles.
@@
Together, these visualizations underscore ReFlow’s ability to maintain sharper physical structures, produce better‐calibrated uncertainties, and faithfully reproduce multiscale spectral statistics.
@@
Figure 19: Richtmyer–Meshkov Mean Field Comparison: Mean velocity fields for the Richtmyer–Meshkov dataset, shown across all physical channels and models. In many cases, predictions closely resemble the ground truth, with differences often imperceptible, underscoring the overall high quality and consistency of the generated mean fields.
@@
Figure 20: Richtmyer–Meshkov Uncertainty Comparison: Per-pixel standard deviation fields from the Richtmyer–Meshkov dataset across all models. Stochastic generative models capture spatially coherent uncertainty patterns, outperforming the deterministic baselines (FNO and UViT), which exhibit limited variability and poor uncertainty localization.
[…]
@@
Figure 22: Richtmyer–Meshkov: spectral and pointwise distribution comparisons between UVit and Rectified Flow.
@@
Figure 23: Shear‐Layer Mean and Standard Deviation Comparison: Visualization of mean-field ($\mu$) and standard-deviation ($\sigma$) predictions for the Shear‐Layer dataset across both velocity components ($u_{x}$, $u_{y}$). For each model, the left half shows the predicted mean fields and the right half shows per-pixel standard deviations. The top row contains the reference ground truth. Despite the reduced channel complexity of this dataset, clear differences emerge in how well models capture coherent structures and localized uncertainties. ReFlow preserves sharper features and more accurately reflects the spatial distribution of uncertainty.
[…]
Figure 25: Shear‐layer: global $u_{x}$ spectra and pointwise $u_{x}$ distributions.
@@
Figure 26: Cloud-Shock Mean Field Comparison: Mean velocity fields for the Cloud-Shock dataset, shown across all physical channels and models. In most cases, predictions closely resemble the ground truth.
@@
Figure 27: Cloud–Shock Uncertainty Comparison: Per-pixel standard deviation visualizations for the Cloud–Shock dataset across all models. While all methods capture key uncertainty patterns, stochastic models produce more localized and physically meaningful variability.
[…]
Figure 29: Cloud‐shock: spectral and pointwise energy comparisons.